OpenStudy (anonymous):

PLEASE HELP!!! Make a table of values, and graph j(x) = 2 x - 2. Describe the asymptote. Tell how the graph is transformed from the graph of f(x) = 2 x.

jimthompson5910 (jim_thompson5910):

is it \[\Large j(x) = \frac{2}{x-2}\] or \[\Large j(x) = \frac{2}{x}-2\] ??

OpenStudy (anonymous):

its j(x)= 2^x-2, sorry i forgot to write it like that

jimthompson5910 (jim_thompson5910):

so \[\Large f(x) = 2^{x}\] and \[\Large j(x) = 2^{x-2}\] ?? or is j(x) \[\Large j(x) = 2^{x}-2\]

OpenStudy (anonymous):

the first one is right, I know that it is transformed 2 units to the right, but i cant find the asymptote

jimthompson5910 (jim_thompson5910):

you are correct, to go from \[\Large f(x) = 2^{x}\] to \[\Large j(x) = 2^{x-2}\] you translate everything 2 units to the right

jimthompson5910 (jim_thompson5910):

the asymptote is a horizontal asymptote this is because as x gets smaller and smaller (in the negative direction), \(\Large 2^{x}\) becomes smaller and smaller (and it essentially becomes 0)

jimthompson5910 (jim_thompson5910):

so the horizontal asymptote for \[\Large f(x) = 2^{x}\] is y = 0

OpenStudy (anonymous):

Thank you so much, my class didn't explain how to find an asymptote or even what it is so it was really confusing, thanks again :)

jimthompson5910 (jim_thompson5910):

the horizontal asymptote for \[\Large j(x) = 2^{x-2}\] is also y = 0

jimthompson5910 (jim_thompson5910):

because shifting everything to the right 2 units will NOT shift the horizontal asymptote